joe is 15 yrs younger then his brother. five years from now, three times his own age will equal twice his brother age how old is joe

Answer by db12345(1)

Basically it is important to convert this into algebra so what I did was:
Let Joe's age = x
Let his brothers age = y
Now we write down all information given algebraically.
x=y-15 (joe is 15 years younger)
in 5 years time --> we add 5 to both sides
x+5=y-10
three times his own age will equal twice his brother age how old is joe --> multiply sides
3(x+5) = 2(y-10)
3x+15=2y-20
3x=2y-5
x=(2y-5)/3 (the "/" means divided by)
solving simultaneously
(2y-5)/3 =y-15 (multiply both sides by 3)
2y-5 = 3y-45
y=40
plug in for x
x= 40-15
x=25
Therefore Joe is 25 years old, feel free to double check by doing
3(25+5) = 90
2(40+5) = 90